A turtle crawls along a straight line, which we will call the x-axis with the positive direction to the right. The equation for the turtle’s position as a function of time is (look at the equation below).


(a) Find the turtle’s initial velocity, initial position, and initial acceleration.

(b) At what time t is the velocity of the turtle zero?

(c) How long after starting does it take the turtle to return to its starting point?

(d) At what times t is the turtle a distance of 10.0 cm from its starting point?


1
Expert's answer
2021-10-31T18:15:41-0400

Assume we have an equation


"x(t)=x_0+bt^2+ct."


(a) Initial position: "x_0". Initial velocity: "c". Initial acceleration: "b".

(b) Determine velocity as the derivative of position by time:


"v(t)=2bt+c,\\
v(t)=0=2bt+c,\\\space\\
t=\frac{-c}{2b}."

(c) The position is zero at:


"bt^2+ct+x_0=0,\\
t_{1,2}=\frac{-c\pm\sqrt{c^2-4bx_0}}{2b}."

(d) Solve with the same equation:

"bt^2+ct+x_0=10,\\
bt^2+ct+(x_0-10)=0,\\
t_{1,2}=\frac{-c\pm\sqrt{c^2-4b(x_0-10)}}{2b}."